OpenStudy (anonymous):

Identify the factors of 4x^2 + 20x + 25. (4x + 5)(x + 5) (4x - 5)(x - 5) (2x + 5)(2x + 5) (2x - 5)(2x - 5) This is A right? @gowtham89

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

Then C?

OpenStudy (anonymous):

Since we have 2 positive signs, let's rule out any of our options with negatives. B and D it can't be. So let's FOIL out A and C

OpenStudy (anonymous):

A gives us a +25 in the middle, so yep, it's C!

OpenStudy (anonymous):

4x^2 + 20x + 25 This fits the form a^2+2ab+b^2 Rewrite it where a=2x and b=5 (2x)^2+2(2x)(5)+(5)^2 2 Apply this rule: a^2+2ab+b^2=(a+b^)2 (2x+5)^2 or (2x+5)(2x+5)

OpenStudy (anonymous):

x is -5/2

OpenStudy (anonymous):

Solve for x over the real numbers: (2 x+5)^2 = 0 Eliminate the exponent. Take the square root of both sides: 2 x+5 = 0 Isolate terms with x to the left hand side. Subtract 5 from both sides: 2 x = -5 Solve for x. Divide both sides by 2: then u ll get x=-5/2

OpenStudy (anonymous):

Identify the factors of 4x^2 + 20x + 25. (4x + 5)(x + 5) (4x - 5)(x - 5) (2x + 5)(2x + 5) (2x - 5)(2x - 5)

OpenStudy (ikram002p):

use FOIL fformula to check which one of them is the factorization \(\large (\overbrace{\overbrace{\color{orange}a +\underbrace{\color{cornflowerblue}b)(\color{seagreen}c}_{\text{Inside}}}^{\text{First}} +\color{brown}d}^{\text{Outside}}) =\overbrace{\color{orange}a\color{seagreen}c}^\text F +\overbrace{\color{orange}a\color{brown}d}^\text O +\underbrace{\color{cornflowerblue}b\color{seagreen}c}_\text I +\underbrace{\color{cornflowerblue}b\color{brown}d}_\text L\\ \qquad\quad \underbrace{\qquad\qquad}_{\text{Last}} \)

OpenStudy (anonymous):

Solve for x over the real numbers: (2 x+5)^2 = 0 Eliminate the exponent. Take the square root of both sides: 2 x+5 = 0 Isolate terms with x to the left hand side. Subtract 5 from both sides: 2 x = -5 Solve for x. Divide both sides by 2: then u ll get x=-5/2

OpenStudy (anonymous):

@BassCatcher15 that is the same question?

OpenStudy (anonymous):

Im sorry. It must not have highlighted. Hang on

OpenStudy (anonymous):

Guys this isn't solving for x this is Identify the factors

OpenStudy (anonymous):

@gowtham89

OpenStudy (anonymous):

Factor completely: -7x^2 + 14x - 21 -1(x^2 - 14x + 21) -7x(x^2 - 2x + 3) -7(x^2 + 2x - 3) -7(x^2 - 2x + 3)

OpenStudy (anonymous):

7 x^2+14 x-21 Factor out the greatest common divisor of the coefficients of -7 x^2+14 x-21. Factor -7 out of -7 x^2+14 x-21: then answer is | -7 (x^2-2 x+3)

OpenStudy (anonymous):

Factor completely: 6x^4y^3 + 21x^3y^2 - 9x^2y 3xy(2x^3y^2 + 7xy - 3x) 3x^2y(2x^2y^2 + 7xy - 3) 3x^2y^3(2x^2 + 7xy - 3) 3x^2y(2x^2y + 7xy - 3y)

OpenStudy (anonymous):

@gowtham89